# Nathalie Gimenes PUC-Rio Tonghui Qi

### v1

@@ 6 Empirical illustration
Our application involves auctions that may proceed to a second round if the property is not sold initially, in which case a new auction is held with a reduced reserve price. In the latter cases, judicial auctions typically apply a 50% discount on the reserve price when a property is not sold in the first round and extrajudicial auctions often set the second-round reserve price to the outstanding debt. To apply our methodology, we always take the reserve price in the first round to be the seller’s optimal reserve price. This is motivated by the fact that the first reserve is set based on actual appraisals of the property value, while the discounted reserve price in the second round is guided by either a judicial mandate or the property debt.
@@
The data we use contains 754 observations covering the period from 2017 to 2023, with the majority corresponding to auctions of type (i) and (ii). We did not collect data for cases of type (iii), as labor courts manage these auctions separately under distinct rules. The sample also excluded auctions deemed as outliers in terms of the reserve price, which we define as having the ratio of the reserve price to the size of the apartment in each auction that is larger than the 99th percentile or smaller than the 1st percentile of the sample. The price of the apartment is in Brazillian reais, and we measure it in R\$$100,000$s. [^10] All prices are expressed in constant January 2017 R\$. We use the size of the apartment (total area measured in sqm) as the covariate, $X$. To compute $W$, i.e. the value of the seller’s outside option, we subtract from the evaluation value of the property the debt registered in its sale report. We only have these information for 341 properties of the total sample. Since we do not use these information to estimate the quantile function of bidder’s private value, the quantiles are estimated using the whole sample. We then use the subsample with evaluation and debt values to estimate the seller’s risk parameter. Throughout this exercise, we assume the seller has a CRRA utility function.
Our estimation procedure then takes place over two stages. First, we estimate the bidder’s private value quantile function for the property. In a second step, we evaluate the seller’s CRRA risk parameter. We estimate $V(\alpha|X)$ using the AQR method for $\alpha=0.01,0.02,...,0.99$. [^11] In a very small proportion of auctions, the observed reserve price is larger than $\widehat{V}(0.99|X)$ or smaller than $\widehat{V}(0.01|X)$. These auctions are removed when estimating $\theta$ and during the model fit analysis. We consider three different bandwidth in our estimation: $h=0.1$, $h=0.15$, and $h=0.2$. The results show that our estimation methodology is robust for different bandwidth choices. To derive the standard error and some quantile levels of $\widehat{\theta}$, we use the nonparametric bootstrap. The bootstrap size is set at 99.
@@ 6.3 Model fit
We see that the distribution of the observed reserve price is almost first order stochastic dominated by the distribution of the counterfactual reserve price when the seller is risk neutral, which implies the risk-neutral seller’s reserve price tend to be higher than the risk-averse counterpart. To provide some quantitative comparisons, we calculate the amount of increase and the percentage increase in the counterfactual reserve price for the whole sample and for three subsamples around each quartile of the apartment size: small group (with area between $20$th and $30$th percentiles), medium group (with area between $45$th and $55$th percentiles), and large group (with area between $70$th and $80$th percentiles). The results are summarized in Table 7.
@@ 7 Conclusion
In this paper we propose a framework to identify and estimate the seller’s risk aversion in ascending auctions. The conditions we use for identification are mild and are analogous to those assumed in the theoretical literature. On the estimation front, we apply the AQR approach of GG22 to estimate valuation quantiles of bidders from the winning bids and to estimate the risk parameter as a two-step semiparametric estimator. Our estimators perform well in a simulation study. We then apply our methodology to study the real estate auction data in Brazil, where we find statistical evidence that the sellers are risk-averse.
While we motivate our interest in seller’s risk aversion with the empirical observation of low reserve prices, it should be mentioned that there are other explanations presented in the theoretical literature to rationalize this. Other possibilities include endogenous entry (Levin and Smith, (1994)), affiliated types (Levin and Smith, (1996)), and risk averse bidders with interdependent values (Hu et al., (2019)). Our paper does not attempt to distinguish between these possibilities. In fact, to date, we are not aware of any works that have attempted this empirically, which would be an interesting direction of research. Nevertheless, we emphasize that allowing sellers to be risk-averse is an important advancement in empirical auctions as some counterfactuals and theoretical results require correct specification of the seller’s risk preference. Currently, empirical studies are assuming sellers are risk-neutral when performing counterfactual studies. Our paper shows how to accommodate risk-averse sellers in an ascending auction.
@@
Since bidders’ behavior in a second-price auction is strategically equivalent to that of an ascending auction under IPV, our estimation strategy in this paper applies to second-price auctions. We believe our approach to model and estimate the seller’s risk aversion can be applied to other types of auctions as well. Beyond the auction format, other interesting aspects of our paper could be explored further. Firstly, we consider the traditional IPV paradigm. Various extensions from this framework have been proposed in the empirical auction literature, such as unobserved heterogeneity (Krasnokutskaya, (2011), Hu et al., (2013), and Luo and Xiao, (2023)), endogenous entry (Marmer et al., (2013), Gentry and Li, (2014), Chen et al., (2025)), and interdependent values (Gimenes and Guerre, (2020)). Even within the narrower domain of AQR applications, which have thus far been applied to ascending and first-price auctions, some practical aspects will benefit from further studies. For instance, while convergence rates for AQR estimators and their corresponding optimal bandwidths are relatively straightforwardly to derive, neither our paper nor GG22 provide practical guidance on bandwidth selection, and both rely on resampling methods for inference without formal justification. Despite encouraging finite-sample performance in Monte Carlo experiments and stable estimates with real data across bandwidth choices, further studies on these aspects will be useful for applied research.
@@
The appendix is organized into three subappendices. Appendix A gives the proof of Proposition 2; Appendix B gives the proofs of results in Section 3; Appendix C gives the proofs of results in Section 4.
@@ B. Asymptotic Theory for Quantile Function Estimators
Before getting into the technical details, it is instructive to clarify how our objects of interest differ from those in GG. In a first-price auction, the quantile function of a bidder’s valuation depends on $B\left(\cdot\right)$ and $B^{\left(1\right)}\left(\cdot\right)$ – the quantile of an individual’s bid and its derivative. GG22 show that the estimation error of $B\left(\cdot\right)$ converges to zero at a faster rate than the counterparts of the estimator of $B^{\left(1\right)}\left(\cdot\right)$, so they derive leading bias and variance expressions for the quantile derivative estimator of the observed bids, but only provide a convergence rate for the quantile estimator itself. In contrast, in an ascending auction, only the quantile of the winning bids is needed to recover the quantile of the bidder’s valuation. Our results provide bias and variance expressions for the quantile estimator in this setting.
Our quantile estimation setup below is a simplification of GG22’s general framework. We use what they call Augmented Quantile Regression (AQR), which is presented in their main text. In contrast, their appendix provides proofs for Augmented Sieve Quantile Regression (ASQR), where the linear quantile specification includes an increasing number of additive terms, leading to a nonparametric specification. As a result, our appendix is much lighter in notation and the proofs simplified.
Our presentation in this sub-appendix aims to be as concise as possible. We outline key arguments, omit steps that follow directly from GG22, refer readers to the original source, and focus only on the elements that require modification from their work. This approach avoids unnecessary repetition of GG’s appendix, which is a diligently crafted piece spanning several sub-appendices that would remain lengthy even without ASQR components.
In what follows, we refer to GG22’s supplementary material as GG-SM. This document can found online at https://doi.org/10.1016/j.jeconom.2021.02.009. The relevant sections are most of Appendices B, C, D, and E. To facilitate readers, we try to use the same notation as in GG22 when possible.
@@ B.1 Outline
In the AQR framework, $\overline{\mathsf{b}}\left(\alpha\right)$ can be interpreted as the pseudo-true parameter. Then $x_{1}^{\top}\mathsf{S}_{0}\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right)$ captures the polynomial approximation bias of $B\left(\alpha|x\right)$. It can be shown that $\widehat{\mathsf{e}}\left(\alpha\right)$ is a zero mean vector, and suitably scaled components of $\widehat{\mathsf{e}}\left(\alpha\right)$ satisfy the central limit theorem (CLT).
@@ B.2 Proofs of results
Next, we consider $J_{S}\left(\alpha,x\right)$. We start with the variance expression for $\widehat{\mathsf{e}}\left(\alpha\right)$ given in GG-SM. In particular, when proving their Lemma B.6 in Appendix F.3.3, they show that $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\left(\mathsf{V}_{e}\left(\alpha\right)+o\left(h\right)\right)/L$, [^13] There are a couple of typos in GG-SM related to $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)$ that are easy to verify. First, at the bottom of their page 134, they were meant to say $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\left(\mathsf{V}_{e}+o\left(h\right)\right)/L$ instead of $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\mathsf{V}_{e}/L+o\left(h\right)$. Second, the last component of $\mathsf{V}_{e}$ given in the second display of page 135 should be $S_{0}^{\top}S_{1}+S_{1}^{\top}S_{0}$ and not $S_{0}S_{1}^{\top}+S_{1}S_{0}^{\top}$. where
@@
It follows that $Var\left(\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)\right)=\alpha\left(1-\alpha\right)\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}/L+O\left(h/L\right)$, and we have $Var\left(\sqrt{L}\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)\right)=\Sigma_{h}\left(\alpha\right)+o\left(1\right)$ as desired.
The central limit theorem (CLT) result follows from the same argument used in the proof of Theorem 3 (and A.3) in GG-SM that can be found in their Appendix E.2. Write $\left(\frac{L}{x_{1}^{\top}\Sigma_{h}\left(\alpha\right)x_{1}}\right)^{1/2}x_{1}^{\top}\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)=\sum\limits_{l=1}^{L}r_{l}\left(\alpha|x\right)$, so that
@@
C. Asymptotic Theory for Risk Aversion Estimator
@@ C.2 Additional Lemmas
Each function in $\mathcal{Q}$ is a sum of three terms. Two of them belong to a low complexity class of functions, $\left\{U_{\theta}:\text{ }\theta\in\Theta\right\}$, that are Lipschitz over $\Theta$ with an $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope. Such a class of functions satisfies the entropy bracketing condition (VW23, Thorem 2.7.17) and is Donsker. The remaining term is a product of three functions: $\left(r,x\right)\mapsto U_{\theta}^{\left(1\right)}\left(r\right)\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)\left(1-\psi_{2}\left(r,x\right)\right)$ where $\left(\psi_{1},\psi_{2}\right)\in\mathcal{D}_{1}\times\mathcal{D}_{2}$. $\left\{U_{\theta}^{\left(1\right)}:\theta\in\Theta\right\}$ is Donsker as it is also a Lipschitz-in-parameter class of functions with an $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope. We have already shown in Lemma C.1 that $\mathcal{D}_{2}$ is Donsker, so that $\left(r,x\right)\mapsto 1-\psi_{2}\left(r,x\right)$ is Donsker. We will show below that $\mathcal{D}_{1\circ 2}=\{$ $\left(r,x\right)\mapsto\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)$ for $\left(\psi_{1},\psi_{2}\right)\in\mathcal{D}_{1}\times\mathcal{D}_{2}$ }\} is also Donsker. Together, this ensures $\mathcal{Q}$ is Donsker as permanence properties of Donsker classes apply to finite sums and products of Donsker classes of functions.
$\mathcal{D}_{1\circ 2}$ is a class of composition of functions with the outer map belonging to $\mathcal{D}_{1}$. The $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope for $\mathcal{D}_{1}$ then applies for functions in $\mathcal{D}_{1\circ 2}$. Note that the standard permanence Donsker theorem employed previously (VW23, Theorem 2.10.8) assumes the outer map of the composition function is fixed, ao it does not apply to $\mathcal{D}_{1\circ 2}$. We need to verify Donsker permanence holds uniformly over $\mathcal{D}_{1}$ as well. To do this, take any $\psi_{1},\psi_{1}^{\prime}$ in $\mathcal{D}_{1}$ and $\psi_{2}\in\mathcal{D}_{2}$, we have

### v2

@@ 6 Empirical illustration
Our application involves auctions that were completed over two rounds. The second round takes place if the property is not sold initially, in which case the reserve price is reduced as judicial auctions typically applied a 50% discount on the reserve price and extrajudicial auctions often set the second-round reserve price to the outstanding debt. To apply our methodology, we take the reserve price in the first round to be the seller’s optimal reserve price. This is motivated by the fact that the first reserve is set based on actual appraisals of the property value, while the discounted reserve price in the second round is guided by either a judicial mandate or the property debt.
@@
The data we use contains 754 observations covering the period from 2017 to 2023, with the majority corresponding to auctions of type (i) and (ii). We did not collect data for cases of type (iii), as labor courts manage these auctions separately under distinct rules. The sample also excluded auctions deemed as outliers in terms of the reserve price, which we define as having the ratio of the reserve price to the size of the apartment in each auction that is larger than the 99th percentile or smaller than the 1st percentile of the sample. The price of the apartment is in Brazillian reais, and we measure it in R\$$100,000$s. [^11] All prices are expressed in constant January 2017 R\$. We use the size of the apartment (total area measured in sqm) as the covariate, $X$. To compute the value of the seller’s outside option, $W$, we subtract from the evaluation value of the property the debt registered in its sale report. We only have these information for 341 properties of the total sample. Since we do not use these information to estimate the quantile function of bidder’s private value, the quantiles are estimated using the whole sample. We then use the subsample with evaluation and debt values to estimate the seller’s risk parameter. Throughout this exercise, we assume the seller has a CRRA utility function.
Our estimation procedure then takes place over two stages. First, we estimate the bidder’s private value quantile function for the property. In a second step, we evaluate the seller’s CRRA risk parameter. We estimate $V(\alpha|X)$ using the AQR method for $\alpha=0.01,0.02,...,0.99$. [^12] In a very small proportion of auctions, the observed reserve price is larger than $\widehat{V}(0.99|X)$ or smaller than $\widehat{V}(0.01|X)$. These auctions are removed when estimating $\theta$ and during the model fit analysis. We consider three different bandwidths in our estimation: $h=0.1$, $h=0.15$, and $h=0.2$. The results show that our estimation methodology is robust for different bandwidth choices. To derive the standard error and some quantile levels of $\widehat{\theta}$, we use the nonparametric bootstrap. The bootstrap size is set at 99.
@@ 6.4 Counterfactual analysis
We see that the distribution of the observed reserve price is almost first order stochastic dominated by the distribution of the counterfactual reserve price when the seller is risk neutral, which implies the risk-neutral seller’s reserve price tend to be higher than the risk-averse counterpart. To provide some quantitative comparisons, we calculate the average of amount increase and the percentage increase in the counterfactual reserve price for the whole sample and for three subsamples around each quartile of the apartment size: small group (with area between $20$th and $30$th percentiles), medium group (with area between $45$th and $55$th percentiles), and large group (with area between $70$th and $80$th percentiles). The results are summarized in Table 7.
@@ 7 Conclusion
In this paper, we propose a framework to identify and estimate the seller’s risk-aversion parameter in ascending auctions. The model is semiparametric, with risk preferences assumed to come from a parametric utility family, while bidder valuations satisfy a linear quantile specification. We show that the risk-aversion parameter is identified under mild conditions that are commonly assumed in the theoretical literature. We then propose a two-step semiparametric estimator for it. This procedure uses the AQR approach of GG22 to estimate bidder valuation quantiles from winning bids in the first stage. We establish that the estimator is consistent and asymptotically normal under standard regularity conditions. Our estimator performs well in a simulation study. We then apply our methodology to real estate auction data in Brazil and find statistical evidence that sellers are risk-averse.
Our study makes a useful contribution to empirical auction research, as optimal auction design can depend crucially on whether the seller is risk averse or risk neutral. Moreover, any supply-side welfare analysis requires correct specification of the seller’s risk preference. Currently, empirical studies assume seller risk neutrality. Learning about seller-side primitives is, however, necessarily more demanding in terms of data requirements than learning about bidder-side primitives, although for the latter—when bid data alone are sometimes sufficient—identification and estimation still depend on the auction format, assumptions about bidders’ information structure, and data availability (e.g., all bids, bids below the reserve, or winning bids). Our application, which uses data from two-stage auction settings, is intended to serve as an illustration of the proposed methodology. More natural settings for its application include auctions with secret or hidden reserve prices, which are common in practice and for which the relevant data are observed by auction platforms and policymakers. While such data are less widely available to researchers, they have been used, for instance, in the first-price auction context (see Elyakime et al., (1994) and Andreyanov and Caoui, (2022)).
It should be noted that seller risk aversion is not the only mechanism that can rationalize low observed reserve prices. Alternative explanations include endogenous entry (Levin and Smith, (1994)), affiliated types (Levin and Smith, (1996)), and risk-averse bidders with interdependent values (Hu et al., (2019)). We do not attempt to distinguish between these possibilities. To our knowledge, no existing empirical work has done so, making this an interesting direction for future studies.
@@
Since bidders’ behavior in a second-price auction is strategically equivalent to that of an ascending auction under IPV, the estimation strategy in this paper applies to second-price auctions. Our general approach to model and estimate the seller’s risk aversion through the revenue maximization condition can be applied to other auction settings. Beyond considering an alternative auction format, the IPV assumption may be relaxed. Various extensions from this framework have been proposed in the empirical auction literature, such as unobserved heterogeneity (Krasnokutskaya, (2011), Hu et al., (2013), and Luo and Xiao, (2023)), endogenous entry (Marmer et al., (2013), Gentry and Li, (2014), Chen et al., (2025)), and interdependent values (Gimenes and Guerre, (2020)). Even within the narrower domain of AQR applications, which have thus far been applied to ascending and first-price auctions, some practical aspects will benefit from further studies. For instance, while convergence rates for AQR estimators and their corresponding optimal bandwidths have been derived, neither our paper nor GG22 provide practical guidance on bandwidth selection, and both rely on resampling methods for inference without formal justification. Despite encouraging finite-sample performance in Monte Carlo experiments and stable estimates with real data across bandwidth choices, further studies on these aspects will be useful for applied research.
@@
The appendix is organized into three subappendices. Appendix A gives the proof of Proposition 2. Appendix B gives the proofs of results in Section 3. Appendix C gives the proofs of results in Section 4.
@@ B. Asymptotic Theory for Quantile Function Estimators
Our quantile estimation setup below is a simplification of GG22’s general framework. We use what they call augmented quantile regression (AQR), which is presented in their main text. In contrast, their appendix provides proofs for augmented sieve quantile regression (ASQR), where the linear quantile specification includes an increasing number of additive terms, leading to a nonparametric specification. As a result, our appendix is lighter in notation and the proofs are simplified.
Our presentation in this sub-appendix aims to be as concise as possible. We outline the key arguments and omit steps that follow directly from GG22. We refer readers to the original source for those steps and focus only on the elements that differ from their work. This approach avoids unnecessary repetition of GG22’s appendix, which is a carefully crafted piece spanning several sub-appendices and would remain lengthy even without the AQR components.
In what follows, we refer to GG22’s supplementary material as GG-SM. This document can found online at https://doi.org/10.1016/j.jeconom.2021.02.009. The relevant sections are most of Appendices B, C, D, and E. To facilitate readers, we use the same notation as in GG22 when possible.
@@ B.1 Outline
In the AQR framework, $\overline{\mathsf{b}}\left(\alpha\right)$ can be interpreted as the pseudo-true parameter. Then $x_{1}^{\top}\mathsf{S}_{0}\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right)$ captures the polynomial approximation bias of $B\left(\alpha|x\right)$. It can be shown that $\widehat{\mathsf{e}}\left(\alpha\right)$ is a zero mean vector, and suitably scaled components of $\widehat{\mathsf{e}}\left(\alpha\right)$ satisfy the CLT.
@@ B.2 Proofs of results
Next, we consider $J_{S}\left(\alpha,x\right)$. We start with the variance expression for $\widehat{\mathsf{e}}\left(\alpha\right)$ given in GG-SM. In particular, when proving their Lemma B.6 in Appendix F.3.3, they show that $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\left(\mathsf{V}_{e}\left(\alpha\right)+o\left(h\right)\right)/L$, [^14] There are a couple of typos in GG-SM related to $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)$. First, at the bottom of their page 134, they were meant to say $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\left(\mathsf{V}_{e}+o\left(h\right)\right)/L$ instead of $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\mathsf{V}_{e}/L+o\left(h\right)$. Second, the last component of $\mathsf{V}_{e}$ given in the second display of page 135 should be $S_{0}^{\top}S_{1}+S_{1}^{\top}S_{0}$ and not $S_{0}S_{1}^{\top}+S_{1}S_{0}^{\top}$. where
@@
It follows that $Var\left(\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)\right)=\alpha\left(1-\alpha\right)\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}/L+O\left(h/L\right)$, and $Var\left(\sqrt{L}\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)\right)=\Sigma_{h}\left(\alpha\right)+o\left(1\right)$ as desired.
The CLT result follows from the same argument used in the proof of Theorem 3 (and A.3) in GG-SM that can be found in their Appendix E.2. Write $\left(\frac{L}{x_{1}^{\top}\Sigma_{h}\left(\alpha\right)x_{1}}\right)^{1/2}x_{1}^{\top}\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)=\sum\limits_{l=1}^{L}r_{l}\left(\alpha|x\right)$, so that
@@
C. Asymptotic Theory for Risk-Aversion Estimator
@@ C.2 Additional Lemmas
An element in $\mathcal{Q}$ is a sum of three terms. Two of them belong to a low complexity class of functions, $\left\{U_{\theta}:\text{ }\theta\in\Theta\right\}$, that are Lipschitz over $\Theta$ with an $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope. Such a class of functions satisfies the entropy bracketing condition (VW23, Theorem 2.7.17) and is Donsker. The remaining (middle) term is a product of three functions: $\left(r,x\right)\mapsto U_{\theta}^{\left(1\right)}\left(r\right)\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)\left(1-\psi_{2}\left(r,x\right)\right)$ where $\left(\psi_{1},\psi_{2}\right)\in\mathcal{D}_{1}\times\mathcal{D}_{2}$. $\left\{U_{\theta}^{\left(1\right)}:\theta\in\Theta\right\}$ is Donsker, as it is also a Lipschitz-in-parameter class of functions with an $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope. We have already shown in Lemma C.1 that $\mathcal{D}_{2}$ is Donsker, so that $\left(r,x\right)\mapsto 1-\psi_{2}\left(r,x\right)$ is Donsker. We will show below that $\mathcal{D}_{1\circ 2}=\{$$\left(r,x\right)\mapsto\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)$ for $\left(\psi_{1},\psi_{2}\right)\in\mathcal{D}_{1}\times\mathcal{D}_{2}$}\} is also Donsker. Together, this ensures $\mathcal{Q}$ is Donsker as permanence properties of Donsker classes apply to finite sums and products of Donsker classes of functions.
$\mathcal{D}_{1\circ 2}$ is a class of composition of functions with the outer map belonging to $\mathcal{D}_{1}$. The $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope for $\mathcal{D}_{1}$ then applies for functions in $\mathcal{D}_{1\circ 2}$. Note that the standard permanence Donsker theorem employed previously (VW23, Theorem 2.10.8) assumes the outer map of the composition function is fixed, so it does not apply to $\mathcal{D}_{1\circ 2}$. We need to verify Donsker permanence holds uniformly over $\mathcal{D}_{1}$ as well. To do this, take any $\psi_{1},\psi_{1}^{\prime}$ in $\mathcal{D}_{1}$ and $\psi_{2}\in\mathcal{D}_{2}$, we have